A buffer solution is prepared by mixing 35mL of 0.18M CH3COOH (acetic acid) and 25mL of 0.23M NaCH3COO (sodium acetate). Find the change in pH when 5.0mL of 0.12M of HCL is added. Please show step by step
Answer #1

pKa of acetic acid = -log[Ka]

                             = -log(1.8 x 10^-5)

                             = 4.74

after addition of HCl = 0.12 M x 5ml = 0.6 mmol

final moles HOAc = 0.18 M x 35 ml + 0.6 = 6.90 mmol

final NaOAc = 0.23 M x 25 ml - 0.6 = 5.15 mmol

using hendersen-hasselbalck equation,

pH = pKa + log(NaOAc/HOAc)

      = 4.74 + log(5.15/6.90)

      = 4.613